Suppose you want to put new carpeting in a 12 ft by 16 ft room. Calculate the amount of carpeting you would need in square yards and calculate the cost of the carpeting if the price is $28.00 per square yard.
Answer by edjones(7311)

12/3*16/3=64/3 yds^2
64/3*28=$597.33
.
Ed